OpenStudy (anonymous):

solve using the addition method 6x+3y=0 8x+5y=8

OpenStudy (anonymous):

How many of these have been done for you this afternoon? Don't you think you could try one on your own?

OpenStudy (anonymous):

none im still working no one give me the answer if could i would that's why we all are here don't you think???

OpenStudy (anonymous):

6x+3y=0 + 8x+5y=8 ------------- 14x+8y=8 Is that the addition method you refer to?